# HiCache: Training-free Acceleration of Diffusion Models via Hermite Polynomial-based Feature Caching

### v1

@@ 4.3 Text-to-Video Generation
Numbers in parentheses show the deviation from the DDIM-50 baseline (FID=2.32, sFID=4.32, IS=241.25).
Table 3: Quantitative comparison on text-to-video generation with HunyuanVideo on Vbench.
@@
Note: All compared methods utilize efficient attention.
Qualitative Analysis. The qualitative results in Figure 8 reveal that competing methods often suffer from temporal inconsistencies, leading to artifacts such as distorted objects and incoherent movements, a common challenge in video generation (Ho et al. 2022b; Singer et al. 2022). In contrast, HiCache maintains high temporal stability and logical consistency, producing more realistic and coherent video sequences.
[…]
4.4 Class-Conditional Image Generation
Quantitative Analysis.
On the task of class-conditional ImageNet generation (Dhariwal and Nichol 2021) with the DiT-XL/2 model (Peebles and Xie 2023), HiCache again demonstrates superior performance. As shown in Table 2, it achieves a lower (better) FID score (Salimans et al. 2016) compared to competing methods at  5x acceleration. This performance gap widens at higher acceleration ratios ( 7.1x), where HiCache maintains a strong FID score while others degrade significantly, further confirming its robustness in high-acceleration scenarios.
4.5 Super-Resolution
On the image super-resolution task with a patch-based Inf-DiT model (Yang et al. 2024), HiCache achieves a significant theoretical ( 5.93×) and wall-clock ( 2.43×) speedup while maintaining near-lossless quality. This task follows the NTIRE 2025 challenge protocol (Chen et al. 2025). As shown in Figure 9, both PSNR and SSIM consistently outperform TaylorSeer, with particularly notable improvements in SSIM. This demonstrates superior stability and quality preservation at high acceleration (detailed results in Appendix A.5).
Figure 9: Performance comparison of HiCache and TaylorSeer on image super-resolution across different acceleration intervals. HiCache consistently outperforms TaylorSeer in PSNR and SSIM, demonstrating superior stability and quality preservation at higher acceleration ratios.
4.6 Ablation Study
To validate the necessity of the contraction factor $\sigma$ introduced in Section 3.3.2, we conduct an ablation study using the same experimental settings as Section 4.2. As shown in Table 4, the results confirm that the contraction mechanism is essential for the method’s viability, with the optimal range representing a careful balance between maintaining the oscillatory properties of Hermite basis functions and ensuring numerical stability, validating our theoretical design that addresses both prediction domain constraints and exponential coefficient growth suppression.
Table 4: Ablation study demonstrating the necessity of the contraction factor $\sigma$ on FLUX for text-to-image generation.
Method FLOPs Speed $\uparrow$ Image Reward $\uparrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
@@ 4.6 Ablation Study
[…]
Note: $\sigma=1.0$ means no contraction factor is applied. Without contraction, HiCache performs worse than TaylorSeer, proving that the dual-scaling mechanism is essential for method viability.
5. Discussion
Empirical Validation of Observation 1
The foundation of HiCache rests on Observation 1, which claims that feature derivative approximations follow a multivariate Gaussian distribution. We validated this using Energy tests (Székely and Rizzo 2005) across 1st to 5th order feature differences in five key FLUX model (Esser et al. 2024) modules. The tests confirmed Gaussianity with maximum confidence (p-value = 1.0) across all 25 configurations, providing strong empirical support for our theoretical framework and justifying Corollary 1 that Hermite polynomials form the optimal basis.
@@ 5. Discussion
To fairly evaluate HiCache against Taylor expansion, we established a simulation framework using real feature trajectories from the FLUX model with activation interval $N_{\text{interval}}=6$, where derivative approximations are computed using sparsely sampled data from previous activation steps. Our simulation eliminates cumulative error by initializing each prediction task with ground-truth historical data. This design isolates the analysis from cascading inaccuracies and enables meaningful quantitative comparison of the basis functions’ intrinsic approximation capabilities, ensuring direct evaluation of performance differences between Hermite and Taylor bases.
@@
Figure 10 provides empirical evidence for the theoretical advantages of Hermite polynomials’ oscillatory properties (discussed in Section 3.3.2). Using the non-cumulative error framework on representative FLUX features, we observe that while both methods perform similarly at order 1, Taylor expansion increasingly over-extrapolates at higher orders, especially at trajectory turning points—precisely confirming Proposition 1. HiCache’s oscillatory Hermite basis, stabilized by the contraction factor $\sigma$ (Definition 2), maintains accurate fitting across all orders.
Quantitative Validation of Proposition 3
@@
To empirically verify the tighter error bounds promised by Proposition 3, we quantitatively compared HiCache against Taylor expansion using the simulation framework. We define the relative error ratio as $R=\nicefrac{{\text{Taylor MSE}}}{{\text{Hermite MSE}}}$ (where $R>1$ indicates superior HiCache performance) and analyzed 25 configurations (5 FLUX modules × 5 polynomial orders).
@@
Figure 11 presents the cumulative error ratios:
Figure 11: Cumulative error ratios (Taylor MSE / HiCache MSE) across 5 FLUX modules as prediction step size varies. HiCache consistently outperforms TaylorSeer (all ratios >1.0>1.0), with the advantage becoming more pronounced at larger prediction steps and higher polynomial orders.
The results empirically confirm HiCache’s consistent advantage ($R>1.0$ across all modules), with the performance gap widening for larger prediction steps and higher polynomial orders. This validates our theoretical analysis: the $O(1/\sqrt{(N+1)!})$ convergence rate from Proposition 3 indeed provides tighter bounds than Taylor’s $O(1/(N+1)!)$, with the advantage becoming more pronounced as $N$ increases—precisely as predicted by Corollary 2.
6. Conclusion
We presented HiCache, a training-free acceleration framework for diffusion models that fundamentally improves feature prediction by replacing Taylor series with scaled Hermite polynomials. Our approach is grounded in rigorous theory: Observation 1 establishes the Gaussian nature of feature derivatives, Corollary 1 identifies Hermite polynomials as the optimal basis, and Proposition 3 provides theoretical guarantees of superior convergence. The scaled Hermite basis (Definition 2) with dual-scaling mechanism $\tilde{H}_{n}(x)=\sigma^{n}H_{n}(\sigma x)$ ensures numerical stability through a single hyperparameter, enabling HiCache to achieve superior performance across diverse tasks:  6.24× speedup on FLUX.1-dev while exceeding baseline quality, maintaining strong FID scores at high acceleration ratios, and pioneering cache-based acceleration for super-resolution. As a plug-and-play solution requiring no model retraining, HiCache establishes a new paradigm for diffusion acceleration by demonstrating that matching basis functions to intrinsic data statistics yields both theoretical and practical benefits.
A.1 Hermite Polynomials: Mathematical Foundations
@@
A.2 Detailed Algorithm Description
@@
5:  $h_{k}\leftarrow\tilde{H}_{k}(\Delta t)$
6:  {This involves computing $H_{k}(\sigma\cdot\Delta t)$ and scaling by $\sigma^{k}$}
7:  $\hat{\mathcal{F}}\leftarrow\hat{\mathcal{F}}+\frac{h_{k}}{k!}\cdot\Delta^{k}\mathcal{F}(t_{last})$ {Accumulate}
8:end for
9:return  $\hat{\mathcal{F}}$
@@
A.3 Complete Error Analysis and Theoretical Guarantees
@@
A.3.1 Truncation Error Analysis
@@
A.3.2 Finite Difference Approximation Error
@@
Under the Gaussian feature dynamics (Observation 1), the expected norm of derivative approximations satisfies:
@@
A.3.3 Numerical Stability Analysis
@@
A.3.4 Total Error Bound and Performance Advantages
@@
$\|\mathbf{E}_{\text{total}}\|$ $\leq\underbrace{C_{2}\frac{(\sigma\sqrt{2}|\Delta s|)^{N_{\text{order}}+1}}{\sqrt{(N_{\text{order}}+1)!}}e^{(\sigma\Delta s)^{2}/2}}_{\text{Truncation}}$
$\quad+\underbrace{C_{4}\Delta t_{\text{hist}}\sqrt{N_{\text{order}}}}_{\text{Approximation}}+\underbrace{C_{5}\epsilon_{\text{machine}}}_{\text{Numerical}}$ (31)
Corollary 3 (Comparative Advantage).
HiCache achieves superior error bounds compared to Taylor expansion through:
  1. 1.
    Faster convergence: $O(1/\sqrt{(N+1)!})$ vs. $O(1/(N+1)!)$
  2. 2.
    Exponential suppression: Factor $\sigma^{N+1}$ with $\sigma<1$
  3. 3.
@@
Proof of Superiority.
The key advantage stems from the denominator $\sqrt{(N+1)!}$ in the Hermite truncation error, which grows much faster than $(N+1)!$ in the Taylor case. Combined with the $\sigma^{N+1}$ suppression factor, this yields exponentially tighter bounds for large $N$. ∎
A.4 Detailed Experimental Setup

### v2

@@ 4.2 Experiment Results
[…]
$\bullet$ denotes the baseline (FLUX.1 [dev] - 50 steps). Best results are in bold and second best are underlined.
$\bullet$ Hi-ClusCa: ClusCa with its Taylor predictor replaced by HiCache’s Hermite predictor.
$\bullet$ FLUX.1 [schnell] is a distilled variant of FLUX.1.
$\bullet$ HiCache+ FLUX.1 [schnell]: Two full computation steps followed by two HiCache-accelerated steps.
Figure 4: Qualitative comparison on the text-to-image task across diverse prompts.
(a) High-frequency detail preservation.
(b) Consistent style and clean backgrounds.
Figure 5: Detail retention and style consistency. (a) Superior detail retention. (b) Greater stability than TaylorSeer under higher acceleration, with consistent style and clean outputs.
Text-to-Video Generation. Table 1 shows HiCache outperforms recent methods. At higher orders ($\mathcal{N}=7$, $\mathcal{O}=2$), HiCache’s advantage over TaylorSeer becomes more pronounced, validating our theoretical analysis. Figure 6 shows HiCache’s superior temporal consistency compared to methods in terms of both temporal consistency and frame generation.
Figure 6: Qualitative comparison of text-to-video generation methods under significant acceleration.
Figure 7: Comparison of HiCache and TaylorSeer on image super-resolution.
Method Latency(s) $\downarrow$ FLOPs(T) $\downarrow$ Speed $\uparrow$ FID $\downarrow$ sFID $\downarrow$ Inception $\uparrow$
@@
$\bullet$ denotes the baseline (DDIM-50 steps). Best results in each block are in bold and second best are underlined.
Class-Conditional Image Generation Table 4 shows that HiCache outperforms recent accelerators on ImageNet with DiT-XL/2 across speedups, with larger relative gains compared with TaylorSeer at higher acceleration—about $3.2\%$ at $\sim 5.0\times$, $4.5\%$ sFID at $\sim 6.2\times$, and $5.9\%/6.0\%$ FID/sFID at $\sim 7.1\times$—while remaining close to the unaccelerated baseline under aggressive acceleration.
Super-Resolution. HiCache achieves $\sim 5.93\times$ theoretical and $\sim 2.43\times$ wall-clock speedup on image super-resolution with Inf-DiT (Yang and others, 2024), following NTIRE 2025 protocol (Chen et al., 2025). Figure 7 shows consistent improvements over TaylorSeer in both PSNR and SSIM. At this acceleration, the restoration quality degrades only mildly: PSNR drops by $\approx 2.24\%$ (from $30.87$ to $30.18$) and SSIM by $\approx 1.44\%$ (from $0.832$ to $0.820$) relative to the interval=1=1 baseline. Detailed latency/FLOPs and restoration metrics for each interval are provided in Table 5.
4.3 Ablation Study
Table 2 dissects our contributions. Without scaling ($\sigma=1.0$), HiCache underperforms TaylorSeer due to numerical instability. With optimal scaling ($\sigma=0.5$), performance improves dramatically. The Hi-Taylor variant (TaylorSeer with HiCache’s dual-scaling) validates that both our scaling mechanism and Hermite basis contribute independently to HiCache’s superiority.
5 Discussion
Empirical Validation of Proposition 2
We validate the Gaussianity of finite differences posited in Proposition 2. Using energy tests (Székely and Rizzo, 2005) across 1st to 5th order feature differences in five key FLUX modules (Esser and others, 2024), the tests confirmed Gaussianity with maximum confidence (p-value = 1.0) across all 25 configurations, providing strong empirical support for our theoretical framework and, in turn, Corollary 1 which motivates Hermite polynomials for Gaussian-correlated processes.
The appendix provides theoretical support through two complementary mechanisms: (i) local linearization yielding conditional Gaussianity for small steps, with residual controlled by second-order bounds, and (ii) aggregation effects invoking multivariate CLT with Berry-Esseen convergence rates. These mechanisms align with transformer architectures’ residual connections and normalization layers that maintain required regularity. Moreover, the robustness analysis shows Hermite bases retain advantages even under approximate Gaussianity, with efficiency degrading at most $O(\epsilon)$ for $\epsilon$-perturbations from exact Gaussian (Appendix G, Proposition 8). This theoretical foundation, combined with empirical energy test validation, justifies modeling finite differences as approximately Gaussian and motivates the Hermite basis choice in Corollary 1.
@@ 5 Discussion
To fairly evaluate HiCache against Taylor expansion, we established a simulation framework using real feature trajectories from the FLUX model with activation interval $N_{\text{interval}}=6$, where derivative approximations are computed using sparsely sampled data from previous activation steps. Our simulation eliminates cumulative error by initializing each prediction task with ground-truth history. This isolates cascading inaccuracies and enables fair quantitative comparison of Hermite and Taylor bases.
@@
Appendix Figure 9 provides empirical evidence for the theoretical advantages of Hermite polynomials’ oscillatory properties (discussed in Section 3.3.2). Using the non-cumulative error framework on representative FLUX features, we observe that while both methods perform similarly at order 1, Taylor expansion increasingly over-extrapolates at higher orders, especially at trajectory turning points—precisely confirming our intuition about monomial limitations. HiCache’s oscillatory Hermite basis, stabilized by the contraction factor $\sigma$ (Definition 2), maintains accurate fitting across all orders (see Appendix H for details).
Quantitative Validation of Error Bounds
@@
To empirically verify the tighter error bounds discussed in our methodology, we quantitatively compared HiCache against Taylor expansion using the simulation framework. We define the relative error ratio as $R=\nicefrac{{\text{Taylor MSE}}}{{\text{Hermite MSE}}}$ (where $R>1$ indicates superior HiCache performance) and analyzed 25 configurations (5 FLUX modules x 5 polynomial orders); Figure 8 summarizes the cumulative error ratios.
@@
Figure 8: Cumulative error ratios (Taylor MSE / HiCache MSE) across 5 FLUX modules.
The results empirically confirm HiCache’s consistent advantage ($R>1.0$ across all modules), with the performance gap widening for larger prediction steps and higher polynomial orders. These results confirm HiCache’s consistent advantage in our parameter regime. In particular, when the contraction condition $\sigma\sqrt{2}\cdot|\Delta s|<1$ holds and for moderate $N$, the empirical truncation behavior aligns with our conditional analysis in Section 3.
6 Conclusion
We introduce HiCache, a training‑free acceleration strategy that overcomes the core limits of Taylor‑series feature caching in diffusion models. Grounded in the empirical observation that feature derivatives follow multivariate Gaussian statistics, we replace monomials with Hermite polynomials—the Karhunen–Loève–optimal basis under Gaussian correlation—and pair them with a simple dual‑scaling mechanism that stabilizes coefficients while retaining the oscillatory behavior needed to model non‑monotonic trajectories. Across text‑to‑image, text‑to‑video, and super‑resolution benchmarks, HiCache consistently matches or surpasses prior methods, delivering large speedups without visible quality loss. Because it only swaps the predictive basis, HiCache drops into existing cache‑then‑forecast pipelines and requires no retraining or architectural changes, offering a principled and practical path to fast diffusion sampling.
Ethics Statement
This work does not introduce new data collection or involve human subjects. All datasets used are publicly available under their original licenses; no personally identifiable or sensitive information is processed. Generated content may reflect biases present in the underlying datasets; we report standard evaluation metrics and release code to facilitate auditing and reproducibility. By reducing the number of expensive model evaluations, HiCache lowers the compute and energy consumption relative to baselines. We discourage misuse of accelerated generative models for deceptive or harmful purposes and will release our implementation with appropriate usage guidelines.
Reproducibility Checklist
1. General Paper Structure
  • $1.$
    Includes a conceptual outline and/or pseudocode description of AI methods introduced (yes/partial/no/NA) yes
  • 1.2.
    Clearly delineates statements that are opinions, hypothesis, and speculation from objective facts and results (yes/no) yes
  • 1.3.
    Provides well-marked pedagogical references for less-familiar readers to gain background necessary to replicate the paper (yes/no) yes
2. Theoretical Contributions
  • 2.1.
    Does this paper make theoretical contributions? (yes/no) yes
    If yes, please address the following points:
    • $2.$
      All assumptions and restrictions are stated clearly and formally (yes/partial/no) yes
    • 2.3.
      All novel claims are stated formally (e.g., in theorem statements) (yes/partial/no) yes
    • 2.4.
      Proofs of all novel claims are included (yes/partial/no) yes
    • 2.5.
      Proof sketches or intuitions are given for complex and/or novel results (yes/partial/no) yes
    • 2.6.
      Appropriate citations to theoretical tools used are given (yes/partial/no) yes
    • 2.7.
      All theoretical claims are demonstrated empirically to hold (yes/partial/no/NA) yes
    • 2.8.
      All experimental code used to eliminate or disprove claims is included (yes/no/NA) yes
3. Dataset Usage
  • 3.1.
    Does this paper rely on one or more datasets? (yes/no) yes
    If yes, please address the following points:
    • 3.2.
      A motivation is given for why the experiments are conducted on the selected datasets (yes/partial/no/NA) yes
    • $3.$
      All novel datasets introduced in this paper are included in a data appendix (yes/partial/no/NA) NA
    • 3.4.
      All novel datasets introduced in this paper will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no/NA) NA
    • 3.5.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are accompanied by appropriate citations (yes/no/NA) yes
    • 3.6.
      All datasets drawn from the existing literature (potentially including authors’ own previously published work) are publicly available (yes/partial/no/NA) yes
    • 3.7.
      All datasets that are not publicly available are described in detail, with explanation why publicly available alternatives are not scientifically satisficing (yes/partial/no/NA) NA
4. Computational Experiments
  • 4.1.
    Does this paper include computational experiments? (yes/no) yes
    If yes, please address the following points:
    • 4.2.
      This paper states the number and range of values tried per (hyper-) parameter during development of the paper, along with the criterion used for selecting the final parameter setting (yes/partial/no/NA) yes
    • 4.3.
      Any code required for pre-processing data is included in the appendix (yes/partial/no) yes
    • $4.$
      All source code required for conducting and analyzing the experiments is included in a code appendix (yes/partial/no) yes
    • 4.5.
      All source code required for conducting and analyzing the experiments will be made publicly available upon publication of the paper with a license that allows free usage for research purposes (yes/partial/no) yes
    • 4.6.
      All source code implementing new methods have comments detailing the implementation, with references to the paper where each step comes from (yes/partial/no) yes
    • 4.7.
      If an algorithm depends on randomness, then the method used for setting seeds is described in a way sufficient to allow replication of results (yes/partial/no/NA) yes
    • 4.8.
      This paper specifies the computing infrastructure used for running experiments (hardware and software), including GPU/CPU models; amount of memory; operating system; names and versions of relevant software libraries and frameworks (yes/partial/no) yes
    • 4.9.
      This paper formally describes evaluation metrics used and explains the motivation for choosing these metrics (yes/partial/no) yes
    • 4.10.
      This paper states the number of algorithm runs used to compute each reported result (yes/no) yes
    • 4.11.
      Analysis of experiments goes beyond single-dimensional summaries of performance (e.g., average; median) to include measures of variation, confidence, or other distributional information (yes/no) yes
    • 4.12.
      The significance of any improvement or decrease in performance is judged using appropriate statistical tests (e.g., Wilcoxon signed-rank) (yes/partial/no) yes
    • 4.13.
      This paper lists all final (hyper-)parameters used for each model/algorithm in the paper’s experiments (yes/partial/no/NA) yes
Appendix A AI Assistance Disclosure
In the preparation of this manuscript, the authors utilized Large Language Models (LLMs) to assist with LaTeX formatting and writing refinement. Specifically, LLMs were employed to help with document structure optimization, mathematical notation consistency, and language polishing to improve clarity and readability. All technical content, experimental results, theoretical contributions, and scientific claims remain entirely the work of the human authors. The use of AI tools was limited to formatting assistance and stylistic improvements, without generating any substantive scientific content or analysis.
Appendix B Hermite Polynomials: Mathematical Foundations
@@
Appendix C Detailed Algorithm Description
@@
5:  $h_{k}\leftarrow\tilde{H}_{k}(-\Delta t)$ {Or equivalently: $h_{k}\leftarrow(-1)^{k}\cdot\tilde{H}_{k}(\Delta t)$}
6:  $\hat{\mathcal{F}}\leftarrow\hat{\mathcal{F}}+\frac{h_{k}}{k!}\cdot\Delta^{k}\mathcal{F}(t_{last})$ {Accumulate}
7:end for
8:return $\hat{\mathcal{F}}$
@@
Appendix D Complete Error Analysis and Theoretical Guarantees
@@
Truncation Error Analysis
@@
Finite Difference Approximation Error
@@
Under the Gaussian feature dynamics assumption, the expected norm of derivative approximations satisfies:
@@
Numerical Stability Analysis
@@
Remark. The uncorrelatedness $\text{Cov}(c_{i},c_{j})=0$ refers to coefficients obtained via $L^{2}(\gamma)$ orthogonal projection; it does not imply statistical independence in general.
@@
Lemma 6 (Envelope bound for Hermite and scaled Hermite).
For all $n\in\mathbb{N}$ and $x\in\mathbb{R}$, there exists a universal constant $C>0$ such that $|H_{n}(x)|\leq C\,e^{x^{2}/2}(\sqrt{2}|x|+1)^{n}\sqrt{n!}$. Consequently, for $\sigma\in(0,1)$, the scaled polynomials satisfy
$|\tilde{H}_{n}(x)|=\sigma^{n}|H_{n}(\sigma x)|\leq C\,e^{(\sigma x)^{2}/2}(\sqrt{2}\,\sigma|x|+1)^{n}\sqrt{n!},$ (25)
exhibiting geometric damping in $n$ controlled by $\sigma$.
Proof.
The bound for $H_{n}$ follows from standard inequalities derived via the generating function
$e^{-t^{2}+2xt}=\sum_{n\geq 0}H_{n}(x)\frac{t^{n}}{n!},$
and Cauchy’s estimates, or from known asymptotics (e.g., Plancherel–Rotach). Multiplying by $\sigma^{n}$ and contracting the input to $\sigma x$ yields the scaled bound. ∎
Total Error Bound and Performance Advantages
@@
$\|\mathbf{E}_{\text{total}}\|\leq\underbrace{C_{2}\frac{(\sigma\sqrt{2}|\Delta s|)^{N_{\text{order}}+1}}{\sqrt{(N_{\text{order}}+1)!}}e^{(\sigma\Delta s)^{2}/2}}_{\text{Truncation}}+\underbrace{C_{4}\Delta t_{\text{hist}}\sqrt{N_{\text{order}}}}_{\text{Approximation}}+\underbrace{C_{5}\epsilon_{\text{machine}}}_{\text{Numerical}}$ (26)
Corollary 2 (Conditional Advantage over Taylor).
Under the condition $\sigma\sqrt{2}\cdot|\Delta s|<1$ and moderate $N$, HiCache can achieve:
  1. 0.1
    Smaller truncation error: The Hermite term $(\sigma\sqrt{2}\cdot|\Delta s|)^{N+1}/\sqrt{(N+1)!}$ can be smaller than Taylor’s $|\Delta s|^{N+1}/(N+1)!$ due to the exponential suppression factor $\sigma^{N+1}$
  2. 0.2
    Oscillatory basis structure: That naturally captures non-monotonic feature dynamics
  3. 0.3
@@
Without the condition $\sigma\sqrt{2}\cdot|\Delta s|<1$, universal faster convergence cannot be guaranteed.
Discussion.
The conditional advantage arises from the interplay between the denominators and the contraction factor. While $1/\sqrt{(N+1)!}$ is actually larger than $1/(N+1)!$, the exponential suppression $\sigma^{N+1}$ (with $\sigma<1$) in the numerator can overcome this difference when $\sigma\sqrt{2}\cdot|\Delta s|<1$, yielding smaller overall error. ∎
Appendix E Detailed Experimental Setup